Historically, the daily average temperature in Waterloo in November has been 2.3 degrees, with a standard deviation of 5 degrees. Assuming daily average temperatures are normally distributed, what is the probability that the mean of the daily average temperatures in Nov 2019 (which has 30 days) will be 5 degrees or more?

Question 8 options:

0.4814

0.0015

0.2946

0.025
Answer #1

Given that ,

mean = \mu = 2.3

standard deviation = \sigma = 5

n = 30

\mu\bar x = 2.3

\sigma\bar x = \sigma / \sqrt n = 5/ \sqrt 30 = 0.9129

P(\bar x >5 ) = 1 - P(\bar x < 5)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (5 - 2.3) /0.9129 ]

= 1 - P(z < 2.96)

Using z table,    

= 1 - 0.9985

= 0.0015

probability = 0.0015

correct option =0.0015